Grade 12Trigonometry

Area of incircle circumcircle and ex circle of a triangle
  • Distance between circuncentre and ortho centre and in centres of atriangle

Varun
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
 
Dear student
 
inradius r = delta /s = area of triangle / semiperimeter
 
also
circum radius R = a/ 2 sinA = b/ 2 sinB = c/ 2 sinC
 
Now you can find area of incircle and circum circle